OpenStudy (anonymous):

Let f(x) = −x2 + 4x − 6. Find and simplify the difference quotient, f(x + h) − f(x) / h.

OpenStudy (gorv):

f(x+h)=-(x+h)^2+4*(x+h)-6=-x^2-h^2-2xh+4x+4h-6) f(x+h)-f(x)=-x^2-h^2-2xh+4x+4h-6-(-x^2+4x-6) f(x+h)-f(x)=-x^2-h^2-2xh+4x+4h-6+x^2-4x+6)=-h^2-2xh+4h=h(-h-2x+4) (f(x+h)-f(x))/h=h(-h-2x+4)/h=-h-2x+4

OpenStudy (gorv):

@pinterx3

OpenStudy (anonymous):

thank you!!

OpenStudy (gorv):

welcome